OpenStudy (anonymous):

factor completely a^4c^4 - 81

OpenStudy (anonymous):

meant a^4x^4- 81

OpenStudy (mertsj):

Do you recognize this as the differrence of 2 squares?

OpenStudy (anonymous):

a^2 - b^2 = (a + b)(a - b)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a^4x^4 - 81 = (a^2x^2 - 9)(a^2x^2 + 9) can it be factored further - i'm looking at the first bracket

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Hint: 9 = 3^2

OpenStudy (mertsj):

Are you with me so far?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Correction: \[a^4x^4-81=[(a^2x^2)^2-9^2]=(a^2x^2-9)(a^2x^2+9)\]

OpenStudy (anonymous):

\[a ^{2}x ^{2} = (ax)^{2} and 9 = 3^{2}\]

OpenStudy (anonymous):

Does that help?

OpenStudy (mertsj):

So now please note that \[a^2x^2-9 \] is also the difference of two squares and can be factored again.

OpenStudy (anonymous):

alright

OpenStudy (mertsj):

\[a^2x^2-9=(ax-3)(ax+3)\]

OpenStudy (mertsj):

Still with me?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

That's it, that's the answer: \[(ax + 3)(ax + 3)(a ^{2}x ^{2} + 9) \]

OpenStudy (mertsj):

So we now have the following for the completely factored form: \[(ax-3)(ax+3)(a^2x^2+9)\]

OpenStudy (anonymous):

Thank you!

OpenStudy (mertsj):

And, of course, the answer pratu posted is incorrect.

OpenStudy (mertsj):

yw. Do you understand?

OpenStudy (anonymous):

can you please give the best answer thing to pratu043 since i cant ?

OpenStudy (anonymous):

Sorry, its ax - 3

OpenStudy (anonymous):

alright & yes i understand thanks!

OpenStudy (anonymous):

You don't need to I gave the wrong answer.

OpenStudy (anonymous):

Thanks anyway.